OpenStudy (anonymous):

help !! white light passes through a diffraction grating of 620 line per mm . First order visible spectrum appear on the wall 50 cm away . If wavelength of white light ranges from 400 nm to 700 nm , how wide is the first order spectrum ?

OpenStudy (anonymous):

@JFraser can you help me with this one ??

OpenStudy (jfraser):

i don't remember this part of spectra, sorry but i'm sure there's an equation somewhere that might help. I like hyperphysics, they do a lot of math, but the diagrams are often very helpful http://hyperphysics.phy-astr.gsu.edu/hbase/phyopt/grating.html

OpenStudy (anonymous):

okay that does help a lot , but , the wavelength is given in range .. what should I do about that ?

OpenStudy (anonymous):

Apply the equation for the largest value and the smallest value (700/400) thus you got the range.

OpenStudy (anonymous):

umm , the range is 1.75 nm ?

OpenStudy (anonymous):

\[d \sin \theta = n \lambda\] d = 10 ^-3/620 ( see that link if you don't know why? http://openstudy.com/study#/updates/53f8cb8ee4b0ae4669f5d505 )

OpenStudy (anonymous):

okay , last answer is 9.3 cm ? right ?

OpenStudy (anonymous):

You will get theta1 and theta 2 at n=1 Then the width = 0.5 (tan theta2 - tan theta1).

OpenStudy (anonymous):

wait if we're using the formula above , what is 50 cm refers to ? is it D or something else ?

OpenStudy (anonymous):

Yea D

OpenStudy (anonymous):

then , why are we using the d sin theta = n lambda

OpenStudy (anonymous):

final answer is 0.11257 meter

OpenStudy (anonymous):

to get theta

OpenStudy (anonymous):

not y = ( m*lambda*D) / d

OpenStudy (anonymous):

no

OpenStudy (anonymous):

OK your question now is what is the equation for width right?

OpenStudy (anonymous):

yes , I'm stuck with the 50 cm and the range wavelength . can't relate those with the question

OpenStudy (anonymous):

 \[\tan \theta = \frac{ y }{ D }\] \[y _{1}= D \tan \theta _{1}.\] \[y _{2}= D \tan \theta _{2}.\] \[width = y _{2} - y _{1} = D \tan \theta _{2} - D \tan \theta _{1} = D ( \tan \theta _{2} - \tan \theta _{1})\]